OpenStudy (scarlettfarra2000):

I need help I'm not so should what I'm supposed to do Simplify the radical expression. Assume that all variables represent positive real numbers

OpenStudy (scarlettfarra2000):

\[5\sqrt{75d ^{2}x}+7d \sqrt{12x}\]

OpenStudy (scarlettfarra2000):

Begin by simplifying each term. The largest square factor of the radicand in the first term. \[\sqrt{75}=15\times5\] \[\sqrt{12}=3\times4\]

OpenStudy (photon336):

@ScarlettFarra2000 so yeah just another thing try to look to see if there are any perfect squares, that are factors of the number you're interested in. let's take your other example for instance . this is what you had here \[\sqrt{12} = \sqrt{4}*\sqrt{3} = 7d*2\sqrt{3}\]

OpenStudy (photon336):

now perfect squares could be 25, 16,4,36, etc

OpenStudy (photon336):

let's take a look at \[\sqrt{75}\] we know that 75 is divisible by 25 right? because \[25*3 = 75\] now check this out it's the same if we put a square root on all of them \[\sqrt{25}*\sqrt{3} =\sqrt{75}\]

OpenStudy (scarlettfarra2000):

Okay, so that's the first step right?

OpenStudy (photon336):

what do you notice about radical 25?

OpenStudy (scarlettfarra2000):

It can b e simplify be 5

OpenStudy (photon336):

so this is the point @ScarlettFarra2000 we need to find a perfect square that is a factor of our number

OpenStudy (photon336):

so take a look at this again \[\sqrt{75} = \sqrt{25}*\sqrt{3} = 5\sqrt{3}\]

OpenStudy (photon336):

\[5*5\sqrt{3d^{2}x}\] now take a look at that d^2x we can take out a d because the square root of d^2 is just d so it becomes \[25d \sqrt{3x}\]

OpenStudy (photon336):

now based on what I did for the first term, how would we simplify the second term?

OpenStudy (scarlettfarra2000):

As we look at \[\sqrt{12}\]\[\sqrt{6}\times \sqrt{2}=\sqrt{12}?\]

OpenStudy (photon336):

one of those numbers has to be a perfect square

OpenStudy (photon336):

@ScarlettFarra2000

OpenStudy (scarlettfarra2000):

yes?

OpenStudy (photon336):

Can you see why it would be this? \[\sqrt{4}*\sqrt{3} = \sqrt{12}\]

OpenStudy (photon336):

what do you notice here that's different from the answer you just gave?

OpenStudy (scarlettfarra2000):

Is it because you can make 4 smaller?

OpenStudy (photon336):

it's because you can simplify 4 to 2 easily

OpenStudy (scarlettfarra2000):

what do you think @robtobey?

OpenStudy (photon336):

@ScarlettFarra2000 what part are you having trouble with?

OpenStudy (scarlettfarra2000):

I'm still not getting it I understand it some like I know how you did some of it like \[25\times5=75 \] and \[3\times4=12\] and I Know why you used those numbers but is that it or is there more to this problem? Or am I over thinking it?

OpenStudy (mathmale):

Given:\[5\sqrt{75d ^{2}x}+7d \sqrt{12x},\]Look for the perfect square factor or factors under each of the radical signs:

OpenStudy (mathmale):

\[5\sqrt{75d ^{2}x}+7d \sqrt{12x}=5\sqrt{3*25d ^{2}x}+7d \sqrt{3*4x}\]

OpenStudy (mathmale):

A factor common to both of the two terms on the right is Sqrt(3x). This can't be simplified further, but the rest of the expression can. Try it.

OpenStudy (sbuck98):

If you have a smartphone, to check your answers download photomath, and type it in and it shows you all the steps.

OpenStudy (scarlettfarra2000):

Thanks @mathmale the answer \[39d \sqrt{3}\]

OpenStudy (mathmale):

Happy to be of help. Photon did a great job of helping also, I thought.

OpenStudy (scarlettfarra2000):

He did I'm just not so great at math

OpenStudy (anonymous):

\[39 d \sqrt{3 x} \]Refer to the attached solution from the Mathematica program.

OpenStudy (scarlettfarra2000):

Thank you @robtobey but I got the answer wrong last time and ever time it's wrong it resets the problem it's almost impossible ever finish this assignment

OpenStudy (mathmale):

Scarlett: It's really, really important that you study this conversation, so that next time you try to work a problem of this kind, you know the proper steps to take. As before, you need to identify any perfect squares that you spot under each radical sign. THAT is the key to simplifying these expressions.

OpenStudy (scarlettfarra2000):

Okay it's just I'm having a REALLY hard time with this lesson some parts I get others not so much